# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v2

@@ I Introduction
Despite the transformative potential of multi-modal diagnostics, several challenges hinder its widespread adoption. The incorporation of molecular data into existing pathology workflows presents substantial complexities, further exacerbating the workload for already overburdened pathologists. Additionally, the scarcity of annotated paired data, due to the resource-intensive and time-consuming nature of labeling, significantly hinders the adoption of supervised learning methods. In this context, multi-modal self-supervised learning (SSL) emerges as a promising alternative, offering the capability to capture robust and comprehensive oncological feature representations without reliance on extensive annotations. Multi-modal SSL methods [8, 9, 10] have demonstrated remarkable success in both natural and medical domains, effectively aligning modalities such as image-text pairs. However, the direct extension of such aligning techniques to histopathology and transcriptomics data presents distinct challenges. Unlike conventional use cases where multi-modal inputs often exhibit highly overlapping features, histopathology and transcriptomics data pairs are inherently more heterogeneous, as they operate at different biological scales and encode distinct yet complementary dimensions of disease-related information. Histopathology provides a morphological and spatial view of tissue architecture, capturing phenotypic traits, while transcriptomics quantifies gene expression levels, uncovering the molecular processes and pathways underlying the disease. Although there are shared correlations between these modalities, each modality also retains substantial modality-specific information [1, 11]. As shown in Figure 1, existing multi-modal methods [12, 13, 14, 15] primarily focus on aligning shared information between modalities, while giving comparatively less attention to the rich and modality-specific information inherent to each data type. For instance, in [15], the authors employed an encoder-only architecture with contrastive learning to enforce representation alignment. However, the optimum training target will be achieved when representations from the same sample become indistinguishable across modalities, thereby eliminating essential modality-specific attributes. Moreover, both histopathology and transcriptomics data contain redundant, disease-unrelated information, including repetitive structural patterns and genes with overlapping functions or pathways. Reducing such redundancies can enable the model to extract more clinically meaningful, disease-relevant representations. Beyond these challenges, the inherent heterogeneity in data formats presents an additional layer of complexity. The histopathology data is structured as 2D image patches, whereas transcriptomic data is represented as tabular numerical values, necessitating careful architectural design.
@@ II-A Self-Supervised Learning in Computer Vision
Recent advancements in SSL in computer vision (CV) have significantly reduced the reliance on labeled data while achieving performance comparable to, or even surpassing, that of supervised methods. SSL can be categorized into three main styles: contrastive learning [18, 19], generative learning [20, 21], and hybrid methods [22, 23] that integrate both approaches.
Multi-modal SSL has further enhanced the capability of models to integrate and align data from different modalities. Among various inputs, vision-language learning [8, 24, 25] is the most extensively studied and demonstrates exceptional performance. Due to the largely overlapping information between image and text data, aligning the two modalities in the latent space alone is often sufficient to achieve impressive results.
@@ II-B Pathological Self-Supervised Learning
To overcome the scarcity of labeled data in Computational Pathology (CPath), a large amount of prior works [26, 27, 28, 29, 30] has explored unsupervised learning and transfer learning techniques. In recent years, SSL methods have gained increasing recognition for their superior potential in CPath, particularly in leveraging large-scale unlabeled Whole Slide Images (WSIs). Similar to general CV, SSL approaches in CPath can be broadly classified into three categories: contrastive [3, 31, 32, 33, 34], generative [35, 36], and hybrid [4, 37, 38, 39].
Multi-modal pathological SSL enhances the model’s ability to learn comprehensive pathological representations by integrating diverse inputs from the diagnostic process. This includes combining WSIs with clinical reports or text captions [40, 41, 42], incorporating transcriptomics data [12], and fusing images from different staining techniques [36, 43]. However, these methods often overlook the inherent redundancy in WSIs and other inputs, leading to the indiscriminate learning of both disease-relevant and disease-irrelevant information, which hampers the models’ performance and efficiency.
@@ II-C Histopathology and Transcriptomics Multi-Modal Learning
In practice, these integration methods predominantly emphasize bridging these modalities through contrastive learning or alternative alignment techniques, such as clustering and cross-attention, to merge data from multiple sources into a unified feature space. Although these modalities share some information, they also contain a vast amount of comprehensive and distinct insights into the disease, reflecting their inherent heterogeneity. While recent advancements have yielded improved outcomes, many existing integration approaches overlook the rich and diverse insights embedded in the unique attributes of each modality.
@@ III Methodology
III-A Modality Encoders
MIRROR adopts two Transformer-based encoders to effectively project histopathology and transcriptomics data into the shared pathological latent space.
@@ III-A2 RNA Encoder
and the model is retrained on the reduced feature set until only $K$ features remain, yielding a compact subset of highly informative genes. Additionally, to further ensure interpretability and biological relevance, manually curated genes associated with specific cancer subtypes are selected based on the COSMIC database [50]. This dual strategy ensures a balance between model performance and biological relevancy as shown in Figure 3.
@@ III-B Preliminaries
where “$\mathrm{r}$” (relevant) vs. “$\mathrm{i}$” (irrelevant) indicates disease relevance, and “$s$” (shared) vs. “$u$” (unique) indicates modality traits. In this notation, the “$+$” denotes the composition of subspace features rather than simple vector addition. While we speak of disease relevance for clarity, in practice these relevant subspaces can capture any shared pathological characteristics that drive meaningful variation in the data. Additionally, for notational conciseness, $\mathbf{S}$ and $\mathbf{S}^{\textit{[CLS]}}$ are treated equivalently in this context.
@@ III-C Modality Alignment Module
The modality alignment module aims to map each encoder output to a common latent space where modality-shared components are systematically brought closer for paired samples, while pusing away irrelevant or mismatched components. To extract the modality-shared information, the encoded representations $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ are mapped into a shared latent space using a modality alignment module for each modality, defined as:
@@
where $B$ is the batch size, and $\tau$ is a temperature hyperparameter.
@@
TABLE I: Cancer subtype classification on TCGA-BRCA, TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
Dataset Backbone Setting ABMIL PORPOISE Linear Classifier TANGLE MIRROR
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
TCGA- NSCLC ResNet-50 10-shot 0.633 0.572 0.742 0.699 0.660 0.658 0.909 0.909 0.930 0.930
±0.084 ±0.147 ±0.182 ±0.242 ±0.024 ±0.025 ±0.033 ±0.033 ±0.006 ±0.006
All Data 0.870 0.868 0.986 0.986 0.845 0.843 0.989 0.987 0.992 0.992
±0.015 ±0.016 ±0.007 ±0.007 ±0.014 ±0.015 ±0.009 ±0.009 ±0.011 ±0.011
Phikon 10-shot 0.709 0.703 0.786 0.767 0.699 0.694 0.925 0.924 0.939 0.938
±0.030 ±0.029 ±0.155 ±0.183 ±0.030 ±0.029 ±0.036 ±0.039 ±0.017 ±0.017
All Data 0.906 0.905 0.984 0.992 0.901 0.900 0.982 0.982 0.994 0.994
±0.010 ±0.010 ±0.007 ±0.008 ±0.008 ±0.008 ±0.012 ±0.012 ±0.007 ±0.007
TCGA- BRCA ResNet-50 10-shot 0.796 0.480 0.902 0.726 0.862 0.507 0.910 0.757 0.910 0.786
±0.075 ±0.028 ±0.010 ±0.152 ±0.031 ±0.076 ±0.036 ±0.121 ±0.032 ±0.092
All Data 0.860 0.462 0.923 0.894 0.901 0.762 0.956 0.909 0.958 0.902
±0.030 ±0.001 ±0.011 ±0.022 ±0.029 ±0.020 ±0.018 ±0.029 ±0.017 ±0.039
Phikon 10-shot 0.712 0.589 0.902 0.731 0.758 0.613 0.921 0.829 0.923 0.832
±0.131 ±0.078 ±0.030 ±0.155 ±0.059 ±0.065 ±0.031 ±0.068 ±0.025 ±0.051
All Data 0.919 0.825 0.934 0.899 0.924 0.821 0.952 0.887 0.955 0.902
±0.008 ±0.021 ±0.030 ±0.009 ±0.038 ±0.092 ±0.022 ±0.063 ±0.023 ±0.047
TCGA- RCC ResNet-50 10-shot 0.752 0.629 0.937 0.846 0.752 0.673 0.938 0.832 0.942 0.853
±0.059 ±0.121 ±0.043 ±0.112 ±0.016 ±0.099 ±0.034 ±0.111 ±0.028 ±0.126
All Data 0.908 0.677 0.976 0.857 0.933 0.827 0.988 0.903 0.998 0.932
±0.017 ±0.144 ±0.031 ±0.194 ±0.007 ±0.116 ±0.007 ±0.138 ±0.003 ±0.150
Phikon 10-shot 0.839 0.733 0.951 0.854 0.828 0.727 0.950 0.867 0.952 0.931
±0.029 ±0.083 ±0.030 ±0.114 ±0.027 ±0.107 ±0.014 ±0.123 ±0.030 ±0.042
All Data 0.968 0.886 0.988 0.920 0.961 0.885 0.989 0.923 0.997 0.931
±0.015 ±0.133 ±0.011 ±0.015 ±0.011 ±0.130 ±0.001 ±0.147 ±0.003 ±0.149
TCGA- COADREAD ResNet-50 10-shot 0.815 0.542 0.813 0.448 0.819 0.502 0.831 0.539 0.834 0.556
±0.057 ±0.066 ±0.065 ±0.020 ±0.055 ±0.080 ±0.043 ±0.066 ±0.039 ±0.144
All Data 0.828 0.515 0.814 0.447 0.816 0.470 0.928 0.881 0.938 0.893
±0.051 ±0.091 ±0.065 ±0.020 ±0.069 ±0.064 ±0.042 ±0.051 ±0.022 ±0.017
Phikon 10-shot 0.659 0.495 0.813 0.448 0.644 0.510 0.772 0.483 0.828 0.515
±0.133 ±0.099 ±0.065 ±0.020 ±0.118 ±0.073 ±0.087 ±0.103 ±0.043 ±0.101
All Data 0.834 0.554 0.859 0.552 0.819 0.570 0.928 0.880 0.941 0.902
±0.037 ±0.114 ±0.065 ±0.217 ±0.069 ±0.096 ±0.049 ±0.046 ±0.020 ±0.018
@@ III-D Modality Retention Module
In contrast to the conventional SSL scenarios where input modalities often exhibit highly overlapping or shared features, histopathology and transcriptomics data are fundamentally heterogeneous. This heterogeneity means that simply aligning these two modalities risks discarding essential modality-specific information encoded in $S_{\mathrm{r},u}^{i},S_{\mathrm{i},u}^{i},T_{\mathrm{r},u}^{i}$, and $T_{\mathrm{i},u}^{i}$. To address this, MIRROR introduces a modality retention module that is explicitly designed to safeguard unique attributes of each modality while allowing the alignment module to focus on shared, cross-modal features.
@@
The reconstructed representation encapsulates both disease-relevant and irrelevant information:
$\mathbf{S}_{\text{retention}}=S_{\mathrm{r},u}+S_{\mathrm{i},u}.$ (11)
@@
Similarly, the retention incorporates both modality-specific and instance-specific information:
$\mathbf{T}_{\text{retention}}=T_{\mathrm{r},u}+T_{\mathrm{i},u}.$ (13)
@@
where $\text{sim}(\cdot)$ denotes the similarity measurement function, for which Mean Square Error (MSE) is used. Given the stochastic nature of masking, where any patch or gene can be masked at random, the encoder is compelled to distribute modality-specific intrinsic structures uniformly across the entire representation. This strategy ensures each modality’s unique biological and morphological characteristics remain well-preserved, improving overall fidelity of the learned representations.
@@ III-E Style Clustering Module
To mitigate redundancy and enhance the representation of disease-relevant information, MIRROR incorporates a style clustering module designed to project both $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ into a shared latent space. Within this module, the latent representations for each sample, denoted as $z_{S}$ and $z_{T}$, are regularized to follow a standard normal distribution $\mathcal{N}(0,I)$. This regularization is enforced by minimizing the Kullback–Leibler (KL) divergence [52]:
@@
$L_{\text{style}}\;=\;\mathrm{KL}\bigl{(}z_{S}\,\big{\|}\,\mathcal{N}(0,I)\bigr{)}\;+\;\mathrm{KL}\bigl{(}z_{T}\,\big{\|}\,\mathcal{N}(0,I)\bigr{)},$ (15)
where the minimization of $L_{\text{style}}$ suppresses unnecessary degrees of freedom in the latent representation, thereby reducing redundancy and promoting a more compact encoding.
To facilitate the alignment of disease-relevant features across modalities, MIRROR introduces a set of learnable cluster centers, $\mathbf{P}\in\mathbb{R}^{D}$, which are normalized to prevent degeneracy. The latent representations $z_{S}$ and $z_{T}$ are softly assigned to these centers, producing assignment distributions $\mathbf{S}_{\text{cluster}}$ and $\mathbf{T}_{\text{cluster}}$, respectively. The alignment between these assignments is encouraged through a bidirectional KL divergence penalty:
$L_{\mathrm{cluster}}\;=\;\mathrm{KL}\bigl{(}\mathbf{S}_{\text{cluster}}\,\|\,\mathbf{T}_{\text{cluster}}\bigr{)}\;+\;\mathrm{KL}\bigl{(}\mathbf{T}_{\text{cluster}}\,\|\,\mathbf{S}_{\text{cluster}}\bigr{)},$ (16)
ensuring mutual consistency between the two modalities.
While both $z_{S}$ and $z_{T}$ are constrained to follow the standard normal distribution $\mathcal{N}(0,I)$, only the disease-relevant consistent pathological signatures of these representations are expected to exhibit meaningful correlations across paired samples. Specifically, pathologically significant phenotypes derived from histopathology data and cancer-related gene expressions or pathways in transcriptomics data are expected to exhibit consistent patterns across modalities. Conversely, variations that are not related to the disease, such as differences in tissue structure that do not correlate with pathological conditions or fluctuations in the expression of non-cancer-related genes, are inherently variable and lack consistency. As a result, these irrelevant factors remain unaligned and do not contribute to minimizing the loss function.
Formally, let $z_{S}$ and $z_{T}$ be $d$-dimensional latent vectors such that:
@@
where $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ encode disease-relevant features, while $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$ capture modality-specific or irrelevant variations. The joint minimization of the style and clustering objectives:
$L_{\text{style clustering}}=L_{\text{style}}+L_{\text{cluster}},$ (18)
encourages the alignment of $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ by fostering similar cluster assignments while suppressing correlations between the irrelevant components $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$.
@@
TABLE II: Cancer survival analysis on TCGA-BRCA and TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
Dataset Backbone Setting ABMIL PORPOISE Linear Classifier TANGLE MIRROR
[…]
TABLE III: Metadata of TCGA cohorts.

### v3

@@ I Introduction
Despite the transformative potential of multi-modal diagnostics, several challenges hinder their widespread adoption. The incorporation of molecular data into existing pathology workflows presents substantial complexities, further exacerbating the workload for already overburdened pathologists. Additionally, the scarcity of annotated paired data, due to the resource-intensive and time-consuming nature of labeling, significantly hinders the adoption of supervised learning methods. In this context, multi-modal self-supervised learning (SSL) emerges as a promising alternative, offering the capability to capture robust and comprehensive oncological feature representations without reliance on extensive annotations. Multi-modal SSL methods [8, 9, 10] have demonstrated remarkable success in both natural and medical domains, effectively aligning modalities such as image-text pairs. However, the direct extension of such aligning techniques to histopathology and transcriptomics data presents distinct challenges. Unlike conventional use cases, where multi-modal inputs often exhibit highly overlapping features, histopathology and transcriptomics data pairs are inherently more heterogeneous, as they operate at different biological scales and encode distinct yet complementary dimensions of disease-related information. Histopathology provides a morphological and spatial view of tissue architecture, capturing phenotypic traits, while transcriptomics quantifies gene expression levels, uncovering the molecular processes and pathways underlying the disease. Although there are shared correlations between these modalities, each modality also retains substantial modality-specific information [1, 11]. As shown in Figure 1, existing multi-modal methods [12, 13, 14, 15] primarily focus on aligning shared information between modalities, while giving comparatively less attention to the rich and modality-specific information inherent to each data type. For instance, in [15], the authors employed an encoder-only architecture with contrastive learning to enforce representation alignment. However, the optimum training target will be achieved when representations from the same sample become indistinguishable across modalities, thereby eliminating essential modality-specific attributes. Moreover, both histopathology and transcriptomics data contain redundant, disease-unrelated information, including repetitive structural patterns and genes with overlapping functions or pathways. Reducing such redundancies can enable the model to extract more clinically meaningful, disease-relevant representations. Beyond these challenges, the inherent heterogeneity in data formats presents an additional layer of complexity. The histopathology data is structured as 2D image patches, whereas transcriptomic data is represented as tabular numerical values, necessitating careful architectural design.
@@ II-A Self-Supervised Learning in Computer Vision
Recent advancements in SSL in computer vision (CV) have significantly reduced the reliance on labeled data while achieving performance comparable to, or even surpassing, that of supervised methods. SSL can be categorized into three main styles: contrastive learning [18, 19], generative learning [20, 21], and hybrid methods [22, 23] that integrate both approaches. Specifically, SimCLR [18] achieves contrastive learning by maximizing agreement between different augmented views of the same image, while MoCo [19] introduces a momentum encoder and a dynamic memory queue to improve stability. BEiT [20] leverages discrete variational autoencoders to tokenize images and reconstruct visual tokens, whereas MAE [21] directly reconstructs raw pixels from masked patches. Hybrid methods such as iBOT [22] combine contrastive and reconstruction on both [CLS] and patch tokens, and DINOv2 [23] incorporates SwAV-style [24] prototype learning with scalable training strategies.
Multi-modal SSL has further enhanced the capability of models to integrate and align data from different modalities. Among various inputs, vision-language learning [8, 25, 26] is the most extensively studied and demonstrates exceptional performance. Due to the largely overlapping information between image and text data, aligning the two modalities in the latent space alone is often sufficient to achieve impressive results. CLIP [8] jointly trains an image encoder and a text encoder to maximize the similarity between paired image-text samples, while minimizing it for unpaired ones. ALIGN [25] extends this training strategy by incorporating noisy image-text pairs, demonstrating the robustness of contrastive alignment at scale. Flamingo [26] builds on this by introducing cross-attention modules that allow the model to condition visual representations directly on textual context, enabling effective few-shot learning across modalities.
@@ II-B Pathological Self-Supervised Learning
To overcome the scarcity of labeled data in Computational Pathology (CPath), a large amount of prior works [27, 28, 29, 30, 31, 32] have explored unsupervised learning and transfer learning techniques. In recent years, SSL methods have gained increasing recognition for their superior potential in CPath, particularly in leveraging large-scale unlabeled Whole Slide Images (WSIs). Similar to general CV, SSL approaches in CPath can be broadly classified into three categories: contrastive [3, 33, 34, 35, 36], generative [37, 38], and hybrid [4, 39, 40, 41]. In the contrastive learning paradigm, HIPT [3] introduces a hierarchical contrastive strategy tailored to multi-resolution WSIs, while CTransPath [33] combines CNN and Transformer backbones to enhance feature generalizability. PathAlign [35] and CPath-Omni [36] extend contrastive SSL to multi-modal contexts by aligning histopathology images with textual descriptions or clinical reports. Among generative methods, MRM [37] proposes masked relation modeling for joint pretraining on medical images and genomics, whereas MMP-MAE [38] leverages mixed attention and modality-specific masked modeling to integrate differently stained images. Hybrid approaches seek to combine the strengths of both paradigms. UNI [4], GigaPath [39], and Virchow [40] adopt DINOv2-based [23] self-distillation frameworks and scale up to millions of pathology slides, aiming to establish general-purpose foundation models. Self-Path [41] employs multi-task self-supervised learning to improve representation quality in low-label regimes.
Multi-modal pathological SSL enhances the model’s ability to learn comprehensive pathological representations by integrating diverse inputs from the diagnostic process. This includes combining WSIs with clinical reports or text captions [42, 43, 44], incorporating transcriptomics data [12], and fusing images from different staining techniques [38, 45]. CONCH [43] leverages the CoCa [46] framework to train on paired WSI-text data, while LLaVA-Med [44] adapts a vision-language assistant for biomedical instruction tuning. MADELEINE [45] improves the model’s ability to interpret differently stained images by jointly pretraining on Hematoxylin and Eosin (H&E) and Immunohistochemistry (IHC) slides. However, these methods often overlook the inherent redundancy in WSIs and other inputs, leading to the indiscriminate learning of both disease-relevant and disease-irrelevant information, which hampers the models’ performance and efficiency.
@@ II-C Histopathology and Transcriptomics Multi-Modal Learning
In practice, these integration methods predominantly emphasize bridging these modalities through contrastive learning or alternative alignment techniques, such as clustering and cross-attention, to merge data from multiple sources into a unified feature space. TANGLE [12] proposed an SSL framework that aligns histopathology slides and transcriptomic profiles through contrastive learning, opening new avenues for multi-modal representation learning in pathology. THREADS [15] extended the scope of data modalities to include DNA, alongside RNA and histology, in a foundation model that still primarily relies on contrastive learning. PIBD [13] employed prototypical clustering and disentanglement strategies to align histology and pathway-level transcriptomics representations. Pathomic Fusion [48] proposed a fusion framework leveraging gated attention and matrix multiplication to integrate WSIs, spatial cell graphs, and genomic features. Zheng et al. [49] applied graph attention networks to combine image and gene expression features, and G-HANet [50] introduced a cross-attention mechanism to dynamically associate histological and genomic representations.
While recent advancements have yielded improved outcomes, it is important to recognize that these modalities, despite sharing some information, also contain a vast amount of comprehensive and distinct insights into the disease, reflecting their inherent heterogeneity. However, current methods often overlook the rich and diverse signals embedded in the unique characteristics of each modality. Xing et al. [51] made a notable advancement by introducing a saliency-aware masking strategy with adaptive distillation, effectively preserving discriminative, modality-specific knowledge and demonstrating the strong potential of masked modeling for nuanced multi-modal representation learning.
@@ III Methodology
III-A Modality Preprocessing and Encoding
MIRROR first applies tailored preprocessing strategies to histopathology and transcriptomics data, including a novel gene selection pipeline for transcriptomics. The processed inputs are then encoded using Transformer-based models to project them into the shared pathological latent space effectively.
@@ III-A2 Transcriptomics Preprocessing
and the model is retrained on the reduced feature set until only $K$ features remain, yielding a compact subset of highly informative genes. Additionally, to further ensure interpretability and biological relevance, manually curated genes associated with specific cancer subtypes are selected based on the COSMIC database [54]. This dual strategy ensures a balance between model performance and biological relevance as shown in Figure 3.
III-A3 RNA Encoder
@@ III-B Preliminaries
where $\mathrm{r}$” (relevant) vs. $\mathrm{i}$” (irrelevant) indicates disease relevance, and “ss” (shared) vs. “uu” (unique) indicates modality traits. In this notation, the “++” denotes the composition of subspace features rather than simple vector addition. While we speak of disease relevance for clarity, in practice, these relevant subspaces can capture any shared pathological characteristics that drive meaningful variation in the data. Additionally, for notational conciseness, $\mathbf{S}$ and $\mathbf{S}^{\textit{[CLS]}}$ are treated equivalently in this context.
@@ III-C Modality Alignment Module
The modality alignment module aims to map each encoder output to a common latent space where modality-shared components are systematically brought closer for paired samples, while pushing away irrelevant or mismatched components. To extract the modality-shared information, the encoded representations $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ are mapped into a shared latent space using a modality alignment module for each modality, defined as:
@@
where $B$ is the batch size, and $\tau$ is a temperature parameter implemented as a learnable scalar, jointly optimized with the model parameters.
@@ III-D Modality Retention Module
In contrast to the conventional SSL scenarios where input modalities often exhibit highly overlapping or shared features, histopathology and transcriptomics data are fundamentally heterogeneous. This heterogeneity means that simply aligning these two modalities risks discarding essential modality-specific information encoded in $S_{\mathrm{r},u}^{i},S_{\mathrm{i},u}^{i},T_{\mathrm{r},u}^{i}$, and $T_{\mathrm{i},u}^{i}$. To address this, MIRROR introduces a modality retention module that is explicitly designed to safeguard unique attributes of each modality without compromising shared semantics.
@@
The reconstructed representation therefore encapsulates modality-specific semantic information while preserving the integrity of shared components:
$\mathbf{S}_{\text{retention}}\approx\mathbf{S}=S_{\mathrm{r},s}+S_{\mathrm{r},u}+S_{\mathrm{i},s}+S_{\mathrm{i},u}.$ (11)
@@
Similarly, the retention incorporates both modality-specific and shared information:
$\mathbf{T}_{\text{retention}}\approx\mathbf{T}=T_{\mathrm{r},s}+T_{\mathrm{r},u}+T_{\mathrm{i},s}+T_{\mathrm{i},u}.$ (13)
@@
where $\text{sim}(\cdot)$ denotes the similarity measurement function, for which Mean Squared Error (MSE) is used. Given the stochastic nature of masking, where any patch or gene can be masked at random, the encoder is compelled to distribute modality-specific intrinsic structures uniformly across the entire representation. This strategy ensures each modality’s unique biological and morphological characteristics remain well-preserved, improving overall fidelity of the learned representations. Because $L_{\text{align}}$ takes effect only on the modality-shared components, the preservation of modality-specific components is guided exclusively by $L_{\text{retention}}$.
@@ III-E Style Clustering Module
To mitigate redundancy and enhance the representation of disease-relevant information, MIRROR incorporates a style clustering module designed to project both $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ into a shared latent space. In this module, each input is first passed through a multilayer perceptron (MLP) to generate a Gaussian distribution in a compressed latent space, parameterized by a mean vector $\mu\in\mathbb{R}^{d}$ and a log standard deviation vector $\log\sigma\in\mathbb{R}^{d}$. A latent code $h\sim\mathcal{N}(\mu,\sigma^{2}I)$ is then sampled using the reparameterization trick [56]. The latent code $h$ is further transformed by a linear projection to produce the final representation $z\in\mathbb{R}^{D}$, where $D$ is the original input dimension. We denote the resulting representations for histopathology and transcriptomics data as $z_{S}$ and $z_{T}$, respectively, and use them for clustering.
To suppress irrelevant or redundant variability, we regularize the latent distributions to remain close to the standard normal prior $\mathcal{N}(0,I)$ by minimizing the Kullback-Leibler (KL) divergence [57]:
@@
$\begin{split}L_{\text{style}}&=\;\mathrm{KL}\bigl{(}\mathcal{N}(\mu_{S},\sigma_{S}^{2}I)\,\big{\|}\,\mathcal{N}(0,I)\bigr{)}\\ &+\,\mathrm{KL}\bigl{(}\mathcal{N}(\mu_{T},\sigma_{T}^{2}I)\,\big{\|}\,\mathcal{N}(0,I)\bigr{)},\end{split}$ (15)
where $(\mu_{S},\sigma_{S})$ and $(\mu_{T},\sigma_{T})$ are the Gaussian parameters estimated from the histopathology and transcriptomics input, respectively. This standard normal prior is the most compact and symmetric non-informative choice under variational approximations, as it introduces no bias toward any specific direction in the latent space. Using a more informative or structured prior could encode unintended signals and weaken the latent compression effect [58, 59]. We prove in Appendix A that Eq. (15) attains a unique global minimum precisely at:
$\mu_{S}=\mu_{T}=0,\quad\sigma_{S}=\sigma_{T}=1.$ (16)
Every bit that drifts away from the prior therefore increases $L_{\text{style}}$, so the network keeps only information that also helps reduce the total loss, effectively pruning disease-irrelevant noise from the latent space.
To facilitate the alignment of disease-relevant features across modalities, MIRROR introduces a set of learnable cluster centers $\mathbf{P}\in\mathbb{R}^{K\times D}$, shared across both modalities. Each cluster center $\mathbf{p}_{k}$ of $\mathbf{P}$ is $\ell_{2}$-normalized to prevent degeneracy. Throughout all experiments, we fix the number of prototypes to K=3, [^0] K=3{,}000, following the default setting in [24, 23]. This prototype-based design promotes alignment by capturing globally consistent, disease-relevant structure rather than enforcing rigid pairwise matching. Unlike hard alignment methods, which can overfit to sample-level noise and lack contextual coherence, the shared cluster centers are learned across the dataset and act as modality-invariant anchors. The use of soft probabilistic assignments further introduces flexibility, allowing each representation to associate with multiple prototypes. This supports uncertainty modeling and avoids over-constraining the latent space, thereby enabling both alignment and retention of clinically meaningful variation.
Given latent representations $z_{S}$ and $z_{T}$, we compute their scaled cosine similarities with the cluster centers $\mathbf{P}$ via a matrix multiplication:
$\mathbf{P}z]_{k}=\langle\mathbf{p}_{k},z\rangle=\|z\|_{2}\cos(\theta_{k}),$ (17)
where $\theta_{k}$ denotes the angle between $z$ and $\mathbf{p}_{k}$. The resulting vector $\mathbf{P}z$ captures the similarity between the representation and each cluster center. These scores are then normalized via softmax to yield probabilistic assignments:
$\mathbf{S}_{\text{cluster}}=\text{softmax}(\mathbf{P}z_{S}),\quad\mathbf{T}_{\text{cluster}}=\text{softmax}(\mathbf{P}z_{T}).$ (18)
These soft assignments represent the affinity of each sample to the $K$ shared cluster centers. Accordingly, $\mathbf{S}_{\text{cluster}},\mathbf{T}_{\text{cluster}}\in\mathbb{R}^{B\times K}$, where each row is a $K$-dimensional probability vector for one sample in the batch. The alignment between these assignments is encouraged through a bidirectional KL divergence penalty:
$L_{\text{cluster}}$ $=\mathrm{KL}\!\bigl{(}\mathbf{S}_{\text{cluster}}\,\|\,\mathbf{T}_{\text{cluster}}\bigr{)}+\mathrm{KL}\!\bigl{(}\mathbf{T}_{\text{cluster}}\,\|\,\mathbf{S}_{\text{cluster}}\bigr{)}$ (19)
$=2\,\mathrm{JSD}\!\bigl{(}\mathbf{S}_{\text{cluster}},\mathbf{T}_{\text{cluster}}\bigr{)}.$
Here, $\mathrm{JSD}(\cdot,\cdot)$ denotes the Jensen–Shannon divergence between two probability distributions. We have proved in Appendix B that $L_{\text{cluster}}$ constitutes an explicit upper bound on the negative mutual information between the style representations, i.e. $I(z_{S};z_{T})\leq g(L_{\text{cluster}})$ for a monotone increasing function $g(\cdot)$. Consequently, minimizing $L_{\text{cluster}}$ increases the mutual information shared by the two modalities.
The total style clustering loss is defined as:
$L_{\text{style clustering}}=L_{\text{style}}+L_{\text{cluster}},$ (20)
and is jointly minimized during training. To analyze how different types of information contribute to the loss, we conceptually decompose the latent vectors $z_{S}$ and $z_{T}$, which are the outputs of the style module, into two parts:
@@
where $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ encode disease-relevant features, while $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$ capture irrelevant variations. This interpretation is used to characterize the distinct roles of latent subspaces during optimization. As shown in Appendix A, $L_{\text{style}}$ penalizes all deviations from the prior, regardless of their semantic content. However, only components that also reduce $L_{\text{cluster}}$ by aligning the soft assignments across modalities contribute to lowering the total loss. This imbalance between the two loss terms naturally encourages the model to retain only informative features. Specifically, the disease-irrelevant components $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$ often reflect factors such as differences in tissue structure that do not correlate with pathological conditions or fluctuations in the expression of non-cancer-related genes. These factors are inherently variable and lack consistency. As a result, they increase $L_{\text{style}}$ but fail to reduce $L_{\text{cluster}}$. Their independence across modalities implies that the prototype assignment distributions derived from these factors contain no mutual information and therefore almost never coincide. Since the Jensen–Shannon divergence in Eq. (19) is equal to zero only when the two distributions are identical, it remains strictly positive and cannot be reduced by modality-specific noise. Consequently, such irrelevant components lead to a higher total loss and are gradually suppressed during training. In contrast, the disease-relevant components $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ may capture features such as pathologically significant phenotypes in histopathology or cancer-associated gene expressions in transcriptomics. Although they may also incur a style penalty, they contribute to reducing $L_{\text{cluster}}$ by improving cross-modal consistency. Since they contribute to minimizing the total loss, these features are preserved.
@@
III-F Global Optimization Objective
The global optimization objective is formally defined as:
TABLE I: Hyperparameter configurations used during pretraining and downstream task evaluations.
Parameter Pretraining Subtyping Survival
10-shot All Data 10-shot All Data
Optimizer Adam Adam Adam
Learning rate $2\times 10^{-5}$ $2\times 10^{-3}$ $3\times 10^{-3}$
Batch size 16 16 16
Epochs 100 100 100
Loss function MIRROR Loss Cross-entropy Loss
Negative Log-Likelihood
Survival Loss
Slide mask ratio 0.75
Transcripts mask ratio 0.75
Mask strategy Random
Initial $\tau$ value $\exp(\log(1/0.07))\approx 14.29$
$L_{\text{align}}$ weight ($\lambda_{\alpha}$) 0.5
$L_{\text{retention}}$ weight ($\lambda_{\beta}$) 0.3
$L_{\text{style clustering}}$ weight ($\lambda_{\gamma}$) 0.2
Number of prototypes 3,000
Frozen encoder Yes Yes
TABLE II: Metadata of TCGA cohorts.